OpenStudy (ajanijones):

Who is good with {Experimental Probability} need help with 5 questions.. please help.

OpenStudy (ajanijones):

say in chat if your good at it.

OpenStudy (ajanijones):

Find the experimental probability. Sheila drew single marbles from a bag, recorded the outcome, and replaced the marble. In 50 draws, Sheila drew 20 white marbles and 30 black marbles. What is the experimental probability that the next marble will be black? A. = 2/5 B. = 3/5 C. = 2/3 D. = 3/7

OpenStudy (e.mccormick):

Experimental Probability is ratio of the number of times an event occurs to the number of tries that have been made. So event:tries aka: event/tries